OpenStudy (australopithecus):

2 lb/ft robe is used to life 800lb of coal up a mine shaft 500ft deep. Find the work done? I need help solving this I have no idea where to start, all I know is I have to use integrals to solve

OpenStudy (australopithecus):

what I know the interval is a=0 b=500 amount of coal = 500lbs weight of rope = 2 lb/ft weight of rope = 1000lbs I'm completely lost

OpenStudy (australopithecus):

I know Work = force * distance and Force = acceleration * mass

OpenStudy (dumbcow):

http://tutorial.math.lamar.edu/Classes/CalcI/Work.aspx it has an example almost exactly like this problem the Force is just the total weight at any given time

OpenStudy (australopithecus):

So I got the solution \[\int\limits_{0}^{500}3(500-x) + 1000dx = 875,000\] what units is this in? also is this answer correct?

OpenStudy (australopithecus):

sorry this is the answer for a question where the coal weighs 1000, and the cable weights 3lbs/ft and the shaft is 500ft deep

OpenStudy (anonymous):

total weight = 2(500) + 800 =1800lb=816.5kg=8165N(taking g=10) displacement=(1/3.281)x500=152m Work done = Weight x displacement=8165 x 152=1241080 J

OpenStudy (dumbcow):

yes answer is correct and units are ft*lbs since the Force is in terms of lbs